Lava Development Company has maintained stable earnings of $7 per share, and has also maintained a 100% payout policy. Recently, Lava has decided to start a new line of business to develop luxury housing with scenic views of active volcanoes. The new line will require Lava to change to a policy of retaining 30% of its earnings for the foreseeable future. You estimate that with the new line of business that dividends will grow by 2% per year in perpetuity (after next year). Lavas current discount rate is 12%. a. Calculate the intrinsic value of Lava Development shares? b. Calculate Lavas net present value of growth opportunities?
Answer #1

Dividend next year = Earnings * ( 1- retention ratio) = 7* (1-30%) = 4.9
Intrinsic Value = Dividend Next Year/( Discount rate - Growth) = 4.9/(12%-2%) = 49
Present Value of Growth Opportunities = Stock Value - Earnings/Discount rate = 49 - 7/12% = 49 - 58.33= -9.33

Best of Luck. God Bless
Please rate WelL